OpenStudy (anonymous):

tampa won 33 fewer than twice as many as they lost. Tampa played 162 games. How many wins, how many losses?

OpenStudy (across):

Tell me, what is confusing you about this problem?

OpenStudy (anonymous):

I have no idea how to complete it...

OpenStudy (across):

Let's break the problem down: Let A be the number of games Tampa won, and let B be the number of games Tampa lost. We know that A + B = 162, and we're told that A = 2B - 33. Do you agree with this? If so, can you solve this system?

OpenStudy (anonymous):

I like how you don't just give the answer.

OpenStudy (anonymous):

me too!

OpenStudy (anonymous):

I agree with it, but I still dont understand...

OpenStudy (anonymous):

how to solve it... I mean do I add 33 to both sides? idk.

OpenStudy (anonymous):

Well, you agree that A + B = 162 is true and A = 2B - 33 is true. Now you have simultaneous equations. You can solve that right?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

A + B = 162 A = 2B - 33 Surely you must have learned about simultaneous equations before this. You have the substitution method and what not.

OpenStudy (anonymous):

I dont understand HOW to do it... I dont know the steps.

OpenStudy (across):

Well, you already know what A is, that is: A = 2B - 33. Why don't you try substituting it in the first equation? E.g. (2B - 33) + B = 162 ?

OpenStudy (across):

Can you solve for B? :)

OpenStudy (anonymous):

65??

OpenStudy (anonymous):

wait..

OpenStudy (anonymous):

21.6?

OpenStudy (anonymous):

the first one was right. How did you get the second one?